- 2/28-3/15: Health check and maintenance for VIS
PRM: One flag was dropped. -> leave it
3/4-14 PRM: Misalignment does not reach where beam hits damper.
3/6 PR3: range -> incline by 20-30mrad at mirror because of bistable.
3/15: BS: Yaw offset adjustment by hand. -> 3/5-6
3/12: BS ok
3/15: some SRs work
3/5 afternoon or 3/6 morning: Coolers for ITMs will be stopped for 2-3 hours to change water.
- 3/22: pump down
3/19: Cooling stop
- 3/28: Opning chamber
- 3/28: Fixing thermal touch
- 3/29: closing chamber and start pumping down.
- 4/1: Start cooling.
1/25-2/20: Yend Pcal finalization
3/5: EYA remaining works by Enomoto, Inoue.
2/25-3/5: Yend WAB finalization
3/6: Y end closing (no laser hazard at least in the morning)
3/7: Y end pump down
- 3/22: leak test
- 3/27: cooling start
2/11-2/14: Xend NAB installation
2/21: Removing FC, Payload release
2/27: ETMX release
2/23-3/6: Xend Pcal finalization 1
3/6: Beam position by IR and Green
3/5 AOS: EY and EX(if they have time) work.
3/6: Green X alignment (no IR), and no hazard at Y end.
3/11, ~20: Xend WAB finalization.
3/11~3/15: ETMX tuning.
3/19: ETMX finalization by Takahashi
3/18,19: ETMX alignement
- 3/23-24: Pcal alignment
- 3/25: Removing PD
- 3/25: closing
- 3/26: Pumping down
- 4/2-4: Leak test
- 3/27: Cooling start
- 3/25~ : Night work officially starts, daily meeting will be moved to 1PM
- Summary of work will last for a while in the morning meeting.
- 2/18-2/22: PRMI commissioning
- 3/25~?: DRMI commissioning
- 3/25-4/19: Y arm commissioning
- 3/25-4/19: FPMI commissioning
- Official night work starts 3/25, but not determined yet.
- Please DO NOT close laser shutter, this is for long term measurement. Currently laser is stopped thin foil at the end of PSL.
- Pcal: (Sat?, Sun) Alignment check at ETMX.
- COM: (Sun) alignment work in the mine.
Report of work
- Type B: It was found that BS was necessary to fix some hight issue yesterday.
- Center area, all the chambers were closed by the end of yesterday, and pump down has just started this morning.
- COM: Alignment was done by the night of day before yesterday.
- Type A:
- ETMX has been confirmed that it is healthy.
- ETMY: IP was something wrong and still alignment was not finished day before yesterday.
- Motor had some problem, and it will be fixed.
- Type B: SR2 oplev is being tuned in these days.
- Drift for all PRMs BS and SRMs will be monitored by Fabian, Terrence and Mark.
- Yend leak test will be done today.